# ViSTA-SLAM: Visual SLAM with Symmetric Two-view Association

### v1

@@ 3.1 Camera Trajectory Evaluation
In Tab. 1 and Tab. 2, we report ATE RMSE. ViSTA-SLAM achieves the best average performance on both datasets, outperforming current SOTA [36] by 17% (0.055 vs. 0.066) and 13% (0.052 vs. 0.060), and surpasses some calibrated methods [61, 25]. ViSTA-SLAM performs less effectively on TUM-RGBD 360 scene due to predominantly rotational camera motion that leads to frontend ambiguity and degrading performance. Other methods [36, 28] use either heavier multi-view frontend or more intensive optimization to reduce the influence. Pure regression-based methods [52, 27] struggle to maintain consistent registration over long sequences with large camera motion due to forgetting effects.
In Fig. 5, we show the estimated trajectories from different methods on 7-Scenes [43] office and TUM-RGBD [45] room. CUT3R [52] suffers from severe forgetting issues on long sequences; SLAM3R [27] has bad point registration on the challenged scene TUM-RGBD room, thus, does not produce correct camera poses. Compared to pure regression-based methods, MASt3R-SLAM [36] and VGGT-SLAM [28] work well, while ViSTA-SLAM achieves even higher trajectory accuracy.
Figure 5: Trajectory estimation results on 7-Scenes office (top) and TUM-RGBD room (bottom). Estimated camera trajectories are projected onto the xxyy plane, with ground-truth shown as dashed lines. The trajectory color encodes ATE RMSE: higher errors in red, lower in blue. For MASt3R-SLAM [36] and VGGT-SLAM [28], only the poses of their selected keyframes are estimated.
@@ 3.2 Dense Reconstruction Evaluation
DROID-SLAM [48] 0.049 0.141 0.048 0.094
MASt3R-SLAM [36] 0.047 0.089 0.085 0.087
Spann3R @20 [50] N/A 0.069 0.047 0.058
Spann3R @2 [50] N/A 0.124 0.043 0.084
CUT3R [52] 0.477 0.276 0.303 0.290
SLAM3R [27] 0.098 0.053 0.059 0.056
MASt3R-SLAM [36] 0.066 0.059 0.056 0.057
VGGT-SLAM [28] 0.070 0.052 0.060 0.056
@@
Table 3: Tracking and Reconstruction Evaluation on 7-Scenes [43]. @$n$ indicates a keyframe every $n$ images. 2-view VGGT w/ PGO uses the 2-view VGGT frontend with the same pose graph optimization as ours. ViSTA-SLAM achieves the best trajectory estimation and reconstruction performance on 7-Scenes.
@@
[52] [27] SLAM [36] SLAM [28] w/ PGO SLAM
@@
Table 4: Model Size and Running Time Evaluation on 7-Scenes [43] redkitchen. Model sizes are reported in billions of parameters. FPS indicates the average number of frames processed per second over three runs. ViSTA-SLAM shows highly competitive real-time speed with the smallest model size among baselines.
In Tab. 4, we evaluate reconstruction quality across methods. Leveraging accurate camera poses and consistent local point clouds, ViSTA-SLAM achieves the best Chamfer distance among all approaches. Despite using a lightweight two-view frontend, ViSTA-SLAM, combined with tailored $\mathrm{Sim}(3)$ pose graph optimization, significantly outperforms multi-view-frontend methods [27, 28] in accuracy (0.45 vs. 0.52) while matching or exceeding completeness. To demonstrate the effectiveness of our lightweight frontend, we add another strong baseline, replacing our STA model with a two-view VGGT [51] as the frontend and conducting the same pose graph optimization. ViSTA-SLAM still achieves better performance in Chamfer distance, completeness, and absolute trajectory error, highlighting the effectiveness of our lightweight symmetric frontend over larger multiview models like VGGT for SLAM tasks
In Fig. 6, we show qualitative reconstruction results on 7-Scenes redkitchen, TUM-RGBD [45] room, and BundleFusion [9] apt1. CUT3R [52] fails to reconstruct correctly due to forgetting issues, while SLAM3R [27] struggles in scenes with large camera perspective changes. MASt3R-SLAM [36] and VGGT-SLAM [28] produce artifacts on object boundaries, failing to clearly separate foreground from background, and show misalignment across views. In contrast, ViSTA-SLAM overcomes these challenges through geometric consistency constraints during training. Notably, VGGT-SLAM fails midway through the apt1 scene as backend optimization diverges, which stems from the unstable RANSAC-based 3D homography estimation, which can sample planar regions and cause ambiguity in their proposed $\mathrm{SL}(4)$ pose graph optimization.
CUT3R [52]
@@
MASt3R-  SLAM [36]
@@
VGGT-  SLAM [28]
@@ 3.3 Model Size and Speed
Table 5: Time Spent on Each Component of ViSTA-SLAM for 7-Scenes [43] redkitchen (in seconds and percentage).
@@
Table 6: Ablation Study on 7-Scenes [43]. w/o pose graph opt. simply accumulates relative poses for absolute poses. w/o two edge types uses the classical pose graph in which each view is represented by a single node.
We compare the frontend model size and processing speed across methods in Tab. 4. Owing to our symmetric design, the decoder and regression heads use only half the parameters of existing feedforward models [53, 23, 50, 27]. Consequently, our model is far more compact: only 64% the size of MASt3R [23] (used in MASt3R-SLAM [36]) and 35% the size of VGGT [51] (used in VGGT-SLAM [28]).
The speed evaluation further confirms that ViSTA-SLAM achieves real-time performance. Benefiting from both the compact frontend and the sparse pose graph, our approach is highly competitive in runtime—faster than the pure regression-based methods CUT3R [52] and SLAM3R [27], and comparable to VGGT-SLAM [28]. It is worth noting that VGGT-SLAM performs inference only once every 32 keyframes, reducing the total number of inference steps. When replacing our STA model with a two-view VGGT that takes two views as input at a time like STA, the running speed is significantly slower, further demonstrating the effectiveness of our lightweight frontend.
@@ 3.4 Ablation Study
Pose graph optimization with loop closure is also highly effective, bringing 48% improvement in the trajectory accuracy (0.105 $\rightarrow$ 0.055), as it introduces simple yet powerful constraints through the edges of the pose graph, preventing error accumulation from two-view estimations as the trajectory grows longer. These findings further support the symmetric formulation of our frontend to regress local pointclouds and relative poses, which maximizes the effectiveness of pose graph optimization since the pointmaps of each view are tightly coupled with their corresponding camera poses in the graph. On the contrary, if pointmaps were regressed in a shared coordinate system inside a submap, as in previous works [53, 23, 50, 5, 27, 52], pose updates would not be able to fix misalignments inside submaps.
@@ 4 Conclusion
{strip}
Method Metric chess fire heads office pumpkin redkitchen stairs Avg
CUT3R Accuracy 0.274 0.102 0.106 0.326 0.452 0.325 0.345 0.276
Completeness 0.303 0.081 0.093 0.441 0.459 0.358 0.389 0.303
Chamfer 0.289 0.091 0.100 0.384 0.456 0.342 0.367 0.290
ATE 0.743 0.226 0.363 0.664 0.546 0.381 0.413 0.477
SLAM3R Accuracy 0.043 0.022 0.020 0.035 0.072 0.062 0.116 0.053
Completeness 0.030 0.013 0.015 0.030 0.055 0.061 0.209 0.059
Chamfer 0.037 0.018 0.017 0.033 0.064 0.061 0.162 0.056
ATE 0.089 0.048 0.036 0.088 0.196 0.102 0.126 0.098
MASt3R-SLAM Accuracy 0.090 0.037 0.027 0.047 0.097 0.070 0.045 0.059
Completeness 0.055 0.024 0.021 0.053 0.054 0.036 0.149 0.056
Chamfer 0.073 0.031 0.024 0.050 0.075 0.053 0.097 0.057
ATE 0.063 0.046 0.029 0.103 0.112 0.074 0.032 0.066
VGGT-SLAM Accuracy 0.029 0.014 0.031 0.041 0.128 0.036 0.087 0.052
Completeness 0.052 0.064 0.021 0.066 0.054 0.057 0.110 0.061
Chamfer 0.040 0.039 0.026 0.054 0.091 0.047 0.098 0.056
ATE 0.037 0.026 0.022 0.103 0.147 0.063 0.095 0.070
Ours Accuracy 0.065 0.015 0.031 0.036 0.061 0.035 0.074 0.045
Completeness 0.063 0.022 0.040 0.048 0.037 0.030 0.154 0.056
Chamfer 0.064 0.019 0.035 0.042 0.049 0.033 0.114 0.051
ATE 0.073 0.035 0.028 0.055 0.129 0.035 0.029 0.055
Table 7: Per Scene Evaluation on 7-Scenes [43]. Comparison of accuracy, completeness, Chamfer distance, and trajectory error on the 7-Scenes dataset. Lower is better. Best results are bold, second best are underlined.
@@ 5 Relative Scale in Pose Graph
6 Per Scene Evaluation Results on 7-Scenes
In Sec. 3.2, only the average reconstruction evaluation is provided. In Tab. 7, we present more detailed per-scene results on 7-Scenes [43] to offer deeper insights. The pure regression method SLAM3R [27] performs well in scenes where the camera primarily focuses on a single corner, such as chess, fire, and heads. However, in scenes involving longer camera trajectories, like pumpkin and redkitchen, its performance degrades due to difficulties in accurately registering points. Our method, ViSTA-SLAM, achieves the best performance in average across all four metrics.
7 More Qualitative Results
@@ 7.1 Pose Graph Optimization
In Fig. 8, we compare the reconstruction and trajectory estimation results with and without pose graph optimization on ScanNet [8] scene0000_00. Pose graph optimization effectively corrects misaligned areas and averages out the errors from the frontend.
@@ 7.3 More Results
   BundleFusion [9] apt2    TUM-RGBD [45] floor
@@
   BundleFusion [9] apt0    7-Scenes [43] pumpkin

### v2

@@ 3.1 Camera Trajectory Evaluation
In Tab. 1 and Tab. 2, we report ATE RMSE. ViSTA-SLAM achieves the best average performance on both datasets, outperforming current SOTA [murai2025mast3rslam] by 17% (0.055 vs. 0.066) and 13% (0.052 vs. 0.060), and surpasses some calibrated methods [zhu2024nicer, lipson2024dpvslam]. ViSTA-SLAM performs less effectively on TUM-RGBD 360 scene due to predominantly rotational camera motion that leads to frontend ambiguity and degrading performance. Other methods [murai2025mast3rslam, maggio2025vggtslam] use either heavier multi-view frontend or more intensive optimization to reduce the influence. Pure regression-based methods [wang2025cut3r, liu2025slam3r] struggle to maintain consistent registration over long sequences with large camera motion due to forgetting effects.
In Fig. 5, we show the estimated trajectories from different methods on 7-Scenes [shotton2013-7scenes] office and TUM-RGBD [sturm12tumrgbd] room. CUT3R [wang2025cut3r] suffers from severe forgetting issues on long sequences; SLAM3R [liu2025slam3r] has bad point registration on the challenged scene TUM-RGBD room, thus, does not produce correct camera poses. Compared to pure regression-based methods, MASt3R-SLAM [murai2025mast3rslam] and VGGT-SLAM [maggio2025vggtslam] work well, while ViSTA-SLAM achieves even higher trajectory accuracy.
Figure 5: Trajectory estimation results on 7-Scenes office (top) and TUM-RGBD room (bottom). Estimated camera trajectories are projected onto the xxyy plane, with ground-truth shown as dashed lines. The trajectory color encodes ATE RMSE: higher errors in red, lower in blue. For MASt3R-SLAM [murai2025mast3rslam] and VGGT-SLAM [maggio2025vggtslam], only the poses of their selected keyframes are estimated.
@@ 3.2 Dense Reconstruction Evaluation
DROID-SLAM [teed2021droid] 0.049 0.141 0.048 0.094
MASt3R-SLAM [murai2025mast3rslam] 0.047 0.089 0.085 0.087
Spann3R @20 [wang2025spann3r] N/A 0.069 0.047 0.058
Spann3R @2 [wang2025spann3r] N/A 0.124 0.043 0.084
CUT3R [wang2025cut3r] 0.477 0.276 0.303 0.290
SLAM3R [liu2025slam3r] 0.098 0.053 0.059 0.056
MASt3R-SLAM [murai2025mast3rslam] 0.066 0.059 0.056 0.057
VGGT-SLAM [maggio2025vggtslam] 0.070 0.052 0.060 0.056
@@
Table 3: Tracking and Reconstruction Evaluation on 7-Scenes [shotton2013-7scenes]. @$n$ indicates a keyframe every $n$ images. 2-view VGGT w/ PGO uses the 2-view VGGT frontend with the same pose graph optimization as ours. ViSTA-SLAM achieves the best trajectory estimation and reconstruction performance on 7-Scenes.
@@
[wang2025cut3r] [liu2025slam3r] SLAM [murai2025mast3rslam] SLAM [maggio2025vggtslam] w/ PGO SLAM
@@
Table 4: Model Size and Running Time Evaluation on 7-Scenes [shotton2013-7scenes] redkitchen. Model sizes are reported in billions of parameters. FPS indicates the average number of frames processed per second over three runs. ViSTA-SLAM shows highly competitive real-time speed with the smallest model size among baselines.
In Tab. 4, we evaluate reconstruction quality across methods. Leveraging accurate camera poses and consistent local point clouds, ViSTA-SLAM achieves the best Chamfer distance among all approaches. Despite using a lightweight two-view frontend, ViSTA-SLAM, combined with tailored $\mathrm{Sim}(3)$ pose graph optimization, significantly outperforms multi-view-frontend methods [liu2025slam3r, maggio2025vggtslam] in accuracy (0.45 vs. 0.52) while matching or exceeding completeness. To demonstrate the effectiveness of our lightweight frontend, we add another strong baseline, replacing our STA model with a two-view VGGT [wang2025vggt] as the frontend and conducting the same pose graph optimization. ViSTA-SLAM still achieves better performance in Chamfer distance, completeness, and absolute trajectory error, highlighting the effectiveness of our lightweight symmetric frontend over larger multiview models like VGGT for SLAM tasks
In Fig. 6, we show qualitative reconstruction results on 7-Scenes redkitchen, TUM-RGBD [sturm12tumrgbd] room, and BundleFusion [dai2017bundlefusion] apt1. CUT3R [wang2025cut3r] fails to reconstruct correctly due to forgetting issues, while SLAM3R [liu2025slam3r] struggles in scenes with large camera perspective changes. MASt3R-SLAM [murai2025mast3rslam] and VGGT-SLAM [maggio2025vggtslam] produce artifacts on object boundaries, failing to clearly separate foreground from background, and show misalignment across views. In contrast, ViSTA-SLAM overcomes these challenges through geometric consistency constraints during training. Notably, VGGT-SLAM fails midway through the apt1 scene as backend optimization diverges, which stems from the unstable RANSAC-based 3D homography estimation, which can sample planar regions and cause ambiguity in their proposed $\mathrm{SL}(4)$ pose graph optimization.
CUT3R [wang2025cut3r]
@@
MASt3R-  SLAM [murai2025mast3rslam]
@@
VGGT-  SLAM [maggio2025vggtslam]
@@ 3.3 Model Size and Speed
Table 5: Time Spent on Each Component of ViSTA-SLAM for 7-Scenes [shotton2013-7scenes] redkitchen (in seconds and percentage).
@@
Table 6: Ablation Study on 7-Scenes [shotton2013-7scenes]. w/o pose graph opt. simply accumulates relative poses for absolute poses. w/o two edge types uses the classical pose graph in which each view is represented by a single node.
We compare the frontend model size and processing speed across methods in Tab. 4. Owing to our symmetric design, the decoder and regression heads use only half the parameters of existing feedforward models [dust3r, mast3r, wang2025spann3r, liu2025slam3r]. Consequently, our model is far more compact: only 64% the size of MASt3R [mast3r] (used in MASt3R-SLAM [murai2025mast3rslam]) and 35% the size of VGGT [wang2025vggt] (used in VGGT-SLAM [maggio2025vggtslam]).
The speed evaluation further confirms that ViSTA-SLAM achieves real-time performance. Benefiting from both the compact frontend and the sparse pose graph, our approach is highly competitive in runtime—faster than the pure regression-based methods CUT3R [wang2025cut3r] and SLAM3R [liu2025slam3r], and comparable to VGGT-SLAM [maggio2025vggtslam]. It is worth noting that VGGT-SLAM performs inference only once every 32 keyframes, reducing the total number of inference steps. When replacing our STA model with a two-view VGGT that takes two views as input at a time like STA, the running speed is significantly slower, further demonstrating the effectiveness of our lightweight frontend.
@@ 3.4 Ablation Study
Pose graph optimization with loop closure is also highly effective, bringing 48% improvement in the trajectory accuracy (0.105 $\rightarrow$ 0.055), as it introduces simple yet powerful constraints through the edges of the pose graph, preventing error accumulation from two-view estimations as the trajectory grows longer. These findings further support the symmetric formulation of our frontend to regress local pointclouds and relative poses, which maximizes the effectiveness of pose graph optimization since the pointmaps of each view are tightly coupled with their corresponding camera poses in the graph. On the contrary, if pointmaps were regressed in a shared coordinate system inside a submap, as in previous works [dust3r, mast3r, wang2025spann3r, cabon2025must3r, liu2025slam3r, wang2025cut3r], pose updates would not be able to fix misalignments inside submaps.
@@ 5 Relative Scale in Pose Graph
6 Additional Quantitative Results
6.1 Per Scene Evaluation Results on 7-Scenes
In Sec. 3.2, only the average reconstruction evaluation is provided. In LABEL:tab:7scenes_results_full, we present more detailed per-scene results on 7-Scenes [shotton2013-7scenes] to offer deeper insights. The pure regression method SLAM3R [liu2025slam3r] performs well in scenes where the camera primarily focuses on a single corner, such as chess, fire, and heads. However, in scenes involving longer camera trajectories, like pumpkin and redkitchen, its performance degrades due to difficulties in accurately registering points. Our method, ViSTA-SLAM, achieves the best performance in average across all four metrics.
6.2 Additional Trajectory Results on TUM-RGBD
In Sec. 3.1, to align with previous methods, only results for the freiburg1 partition of the TUM RGB-D dataset [sturm12tumrgbd] are reported. In Tab. 8, we also report results for the freiburg2 and freiburg3 partitions.
Sequence ATE RMSE (m)
freiburg2_360_hemisphere 0.2037
freiburg2_360_kidnap 0.4617
freiburg2_desk 0.0577
freiburg2_large_with_loop 0.2170
freiburg2_rpy 0.0222
freiburg2_xyz 0.0155
freiburg3_cabinet 0.3869
freiburg3_large_cabinet 0.1334
freiburg3_long_office_household 0.1013
freiburg3_teddy 0.0789
Table 8: Trajectory ATE results on the freiburg2 and freiburg3 partitions of the TUM RGB-D dataset [sturm12tumrgbd].
6.3 Additional Evaluation on More Datasets
In Tab. 10 and Tab. 10, we additionally report camera trajectory evaluation results on Replica [replica19arxiv] and ScanNet [dai2017scannet], as well as reconstruction evaluation results on Replica for several commonly used SLAM testing scenes.
Scene ATE Acc. Comp. Chamfer
office0 0.0744 0.0595 0.0226 0.0410
office1 0.1934 0.2614 0.1833 0.2223
office2 0.1177 0.0914 0.0316 0.0615
office3 0.0485 0.0623 0.0221 0.0422
office4 0.1302 0.1338 0.0688 0.1013
room0 0.0688 0.0766 0.0209 0.0488
room1 0.0934 0.1105 0.0552 0.0828
room2 0.1363 0.1194 0.0223 0.0709
Table 9: Per-scene evaluation results on Replica [replica19arxiv].
   Sequence    ATE RMSE (m)
   scene0000_00    0.0483
   scene0059_00    0.0391
   scene0106_00    0.0559
   scene0169_00    0.0526
   scene0181_00    0.0520
   scene0207_00    0.0479
Table 10: Per-scene camera trajectory evaluation results on ScanNet [dai2017scannet].
@@ 6.3 Additional Evaluation on More Datasets
7 Additional Qualitative Results
@@ 7.1 Pose Graph Optimization
In Fig. 8, we compare the reconstruction and trajectory estimation results with and without pose graph optimization on ScanNet [dai2017scannet] scene0000_00. Pose graph optimization effectively corrects misaligned areas and averages out the errors from the frontend.
@@ 7.3 More Results
   BundleFusion [dai2017bundlefusion] apt2    TUM-RGBD [sturm12tumrgbd] floor
@@
   BundleFusion [dai2017bundlefusion] apt0    7-Scenes [shotton2013-7scenes] pumpkin